OpenStudy (henrietepurina):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

OpenStudy (henrietepurina):

@adrynicoleb @Awesome781 @AlexaByers98 @arilove1d @ams98

OpenStudy (henrietepurina):

help someone

OpenStudy (ranga):

Slope is -6. That quickly eliminates 2 and 4 because the slope is +6. Try x = -2 and y = -3 in 1 and 3 and see which one is true.

OpenStudy (henrietepurina):

i need it by 9:50 just tell me for a medal cuz i have more to go

OpenStudy (ranga):

Sorry. I am not allowed to give out answers but only help students learn.

OpenStudy (henrietepurina):

ok well what should I do first?

OpenStudy (ranga):

Try the first.

OpenStudy (henrietepurina):

y = -6(x + 2) - 3 y = -6x -12 - 3 y = -6x -15

OpenStudy (ranga):

I will tell you another method which might be faster. A line passing through (x1, y1) with a slope m has the equation: y - y1 = m(x - x1) Here the point is (-2, -3). So x1 = -2, y1 = -3 and slope m = 6 Just plug in the numbers into (1). No need to simplify. Then check the answers.

OpenStudy (ranga):

Just plug in the numbers into y - y1 = m(x - x1)

OpenStudy (henrietepurina):

ok so y - (-3) = 6( x - (-2))

OpenStudy (ranga):

sorry m = -6.

OpenStudy (ranga):

Yes. y - (-3) = -6(x - (-2)) y + 3 = -6(x + 2)

OpenStudy (henrietepurina):

what next?

OpenStudy (ranga):

There is your answer. Check which choice is y + 3 = -6(x + 2)

OpenStudy (henrietepurina):

there is none like that

OpenStudy (henrietepurina):

maybe y = -6x + 2 - 3 y = -6x -1

OpenStudy (ranga):

Don't you think the answer y + 3 = -6(x + 2) matches the first choice?

OpenStudy (henrietepurina):

oh it does? thanks!

OpenStudy (ranga):

yw.